### v3

Abstract
While computer vision has achieved tremendous success with multimodal encoding and direct textual interaction with images via chat-based large language models, similar advancements in medical imaging AI—particularly in 3D imaging—have been limited due to the scarcity of comprehensive datasets. To address this critical gap, we introduce CT-RATE, the first dataset that pairs 3D medical images with corresponding textual reports. CT-RATE comprises 25,692 non-contrast 3D chest CT scans from 21,304 unique patients. Through various reconstructions, these scans are expanded to 50,188 volumes, totaling over 14.3 million 2D slices. Each scan is accompanied by its corresponding radiology report. Leveraging CT-RATE, we develop CT-CLIP, a CT-focused contrastive language-image pretraining framework designed for broad applications without the need for task-specific training. We demonstrate how CT-CLIP can be used in two tasks: multi-abnormality detection and case retrieval. Remarkably, in multi-abnormality detection, CT-CLIP outperforms state-of-the-art fully supervised models across all key metrics, effectively eliminating the need for manual annotation. In case retrieval, it efficiently retrieves relevant cases using either image or textual queries, thereby enhancing knowledge dissemination. By combining CT-CLIP’s vision encoder with a pretrained large language model, we create CT-CHAT, a vision-language foundational chat model for 3D chest CT volumes. Finetuned on over 2.7 million question-answer pairs derived from the CT-RATE dataset, CT-CHAT surpasses other multimodal AI assistants, underscoring the necessity for specialized methods in 3D medical imaging. Collectively, the open-source release of CT-RATE, CT-CLIP, and CT-CHAT not only addresses critical challenges in 3D medical imaging, but also lays the groundwork for future innovations in medical AI and improved patient care.
Introduction
Three-dimensional (3D) medical imaging, such as computed tomography (CT) and magnetic resonance imaging (MRI), provides a more detailed and comprehensive view of a patient’s condition compared to two-dimensional (2D) imaging [1]. Among these, CT stands out due to its speed and widespread availability, making it the preferred modality for diagnosing various abnormalities [2]. As global demand for CT scans continues to grow, the need for rapid and accurate diagnostic results becomes more pressing. However, the manual generation of reports from these complex images is time-consuming and prone to errors, underscoring the urgent need for automation [3]. Artificial intelligence (AI) has the potential to address this challenge by efficiently identifying pathologies, reducing the workload on radiologists, and potentially saving lives in situations where immediate radiological expertise is not available.
AI has already begun to revolutionize the interpretation of medical images [4, 5, 6, 7, 8]. Recent advances in automated diagnosis, classification, and segmentation have been driven by numerous publicly available datasets [9, 10, 11, 12, 13]. Beyond supervised methods, general AI systems have been developed using open-source image-text pairs in self-supervised learning, particularly in 2D imaging [14]. These versatile AI foundation models integrate comprehensive natural language annotations into the learning process without requiring task-specific training, thereby eliminating the need for explicit manual image annotations. Unconstrained by predefined classes or tasks during training, they enhance adaptability and have been successfully applied to tasks such as disease classification [15], case retrieval [16], and interactive diagnostics via chat interfaces [14]. Despite these significant advancements in 2D imaging, the application of advanced AI techniques to 3D imaging remains underexplored. A major challenge is the lack of open-source 3D medical imaging datasets paired with textual reports, which require extensive resources, specialized knowledge, and technical expertise to create [17, 18, 19]. Currently, the largest 3D medical imaging datasets available contain only a few thousand cases and do not include paired textual data [20, 21, 22], limiting the development of advanced AI models. Moreover, 3D models face higher computational demands and have fewer available pretrained open-source models compared to their 2D counterparts [23, 24]. Previous studies have only begun to address these challenges; for example, one involving chest CT is limited by its supervised approach, restricting detection to predefined abnormalities and offering only a small portion of its dataset publicly without paired textual reports [20].
To address these challenges, we present three related contributions: First, we introduce CT-RATE, a dataset that pairs 50,188 reconstructed 3D chest CT volumes (over 14.3 million 2D slices) from 25,692 scans of 21,304 patients with corresponding radiology reports. Each volume is annotated with 18 abnormalities, supporting extensive validation and fine-tuning. CT-RATE is fully open-source, making it the first public 3D medical imaging dataset with paired textual reports and the largest CT dataset with multi-abnormality labels. Second, leveraging CT-RATE, we develop CT-CLIP, a novel chest CT-focused contrastive language-image pretraining framework. CT-CLIP integrates extensive natural language annotations into the learning process, enabling the model to comprehend image-based semantic knowledge and perform various tasks without specific training. Our evaluation shows CT-CLIP’s ability to detect multiple abnormalities in a zero-shot setting [25], outperforming state-of-the-art fully supervised methods on three test sets from two different countries. These findings indicate that explicit labels are not essential for achieving high performance in 3D medical imaging interpretation, given the availability of corresponding radiology reports. Third, to extend CT-CLIP’s capabilities, we combine its vision encoder with a pretrained large language model to create CT-CHAT, a vision-language chat model for 3D chest CT volumes. CT-CHAT is fine-tuned with over 2.7 million question-answer pairs derived from the CT-RATE dataset. CT-CHAT significantly outperforms open-source conversational vision-language AI assistants, such as LLaVA 1.6 [26], LLaVA-Med [27], and CXR-LLaVA [28], underscoring the need for specialized tools in 3D medical imaging. We also benchmark CT-CHAT against state-of-the-art 3D medical report generation methods, including RadFM [29] and CT2Rep [30], using clinically meaningful accuracy metrics. On two independent test sets, CT-CHAT outperforms these baselines by a significant margin, demonstrating both its clinical relevance and technical efficacy in report generation.
In this article, we introduce pivotal contributions to medical AI. We address the scarcity of open-source 3D imaging datasets paired with textual reports by introducing CT-RATE. Then, we create CT-CLIP, a versatile foundation model that effectively integrates linguistic and visual data from CT-RATE via contrastive learning [31]. We demonstrate CT-CLIP’s remarkable adaptability and performance across a variety of tasks without task-specific training. Moreover, we develop a vision-language AI assistant, CT-CHAT, by finetuning a pretrained large language model through visual instruction using CT encodings created by CT-CLIP’s vision encoder. The open-source release of CT-RATE, CT-CLIP, and CT-CHAT addresses critical challenges in 3D medical imaging, such as dataset scarcity, model versatility, and learning paradigms, leading to groundbreaking research and substantial advancements in the field.
Results
CT-RATE: Pairing CT volumes with radiology text reports
We curate and open-source CT-RATE, a dataset comprising 25,692 non-contrast 3D chest CT volumes and corresponding radiology text reports from 21,304 unique patients (Figure 1a). These 3D volumes, reconstructed using various methods to suit different window settings [32], total over 14.3 million 2D slices. Each CT exam is accompanied by a report that includes an impression and a findings section, authored by a radiologist, along with additional sections detailing patient information and scan technique. CT-RATE is divided into training and internal validation sets for experiments (see creating the CT-RATE dataset section of Methods). While CT-RATE is a single-center dataset, it captures substantial heterogeneity in CT scanners, acquisition protocols, and reconstruction techniques. To evaluate the generalization of our models (developed on CT-RATE) under distribution shifts, we incorporate two external validation sets from independent studies conducted in different countries [20, 33] (see adding external validation datasets to CT-RATE section of Methods). First, we utilize the public RAD-ChestCT dataset, which includes multi-abnormality labels, to assess model generalizability. Second, we perform a real-world validation by inviting an independent clinical team—uninvolved in model development—to download our open-source codebase, models, and user interface from the public repository and apply them to their own local, private dataset without any adaptation. The performance scores they obtain are presented in the following sections. Notably, none of the external validation data are used during development; all model development is conducted exclusively using CT-RATE.
Conventional supervised machine learning approaches for medical image classification necessitate manual annotation of large datasets, a process that requires significant time and expertise [34]. Although our self-supervised learning approach does not require abnormality labels, the baseline supervised model and finetuning methods do. To meet this need, we enrich CT-RATE with global abnormality labels extracted from the radiology reports using an automated text classifier (Figure 1b). This process involves finetuning a pretrained language model, RadBERT [35], similar to recent works [36]. Within CT-RATE, we manually annotate 1,000 reports for 18 distinct abnormalities for this finetuning procedure. The finetuned model’s precision, recall, and F1 scores, detailed in Supplementary Table 2, show robust classification across abnormalities, ensuring the CT-RATE dataset is enriched with high-quality labels.
Figure 2 summarizes the CT-RATE dataset alongside our open-source external validation set, RAD-ChestCT. Our dataset spans a wide age range (18-102 years) (Figure 2a) and includes diverse pathologies primarily related to both heart and lung conditions (Figure 2b). The reports vary in length and are authored by numerous radiologists, reflecting a variety of writing styles (Figure 2c). The dataset also varies in CT slice numbers (Figure 2d), slice thickness, axial pixel (XY) spacing (Figure 2e), and includes data from different CT manufacturers and a balanced sex distribution (Figure 2f). This heterogeneity is crucial for developing robust models capable of handling out-of-distribution cases. Supplementary Table 1 provides detailed statistics on abnormalities, manufacturers, and sex distributions in the training and internal validation sets. In conclusion, CT-RATE does not represent a curated study dataset, but rather reflects real-world scenarios characterized by heterogeneity. This diversity ensures that the models trained on CT-RATE are well-suited for deployment in clinical settings, offering reliable performance across a broad range of conditions and imaging scenarios.
CT-CLIP: Developing a visual-language foundation model with CT-RATE
Unlike other methods utilizing supervised machine learning for multi-abnormality classification of CT volumes, our study employs radiology reports, which provide a rich source of semantic knowledge. This significantly enhances the understanding of CT volumes and supports a wide range of potential zero-shot applications. Accordingly, we develop CT-CLIP, the first foundational model for chest CT-focused contrastive language-image pretraining, adapting a technique from a previous study [37]. We utilize both the impression and findings sections of radiology reports combined for CT-CLIP due to their superior performance compared to utilizing either one alone, as demonstrated by the abnormality classification scores in an ablation study (Supplementary Figure 1a). While the best performance is achieved training on both impression and findings together, using only the findings section performs better than using only the impression section, indicating that the findings section contains more clinically informative content.
CT-CLIP comprises a vision transformer and a text transformer that generate two distinct embedding vectors: one for chest CT volumes and another for reports, respectively (see developing the CT-CLIP model section of Methods). The vision transformer is adapted from CT-ViT [23], incorporating a two-stage spatial and causal transformer, enabling 3D attention across volumes. CT-CLIP is trained to maximize the similarity between embedding vectors for paired volume-text data samples, while minimizing similarity for unpaired data. Our training approach leverages contrastive learning [38] (Figures 1c, 1d), which enables the training of a foundation model suitable for various applications without task-specific training, including multi-abnormality detection, volume-to-volume retrieval, and report-to-volume retrieval. Subsequent sections demonstrate how CT-CLIP can be applied to these downstream tasks without additional training. To further enhance CT-CLIP’s capabilities, we explore finetuning methods for detection to achieve greater accuracy and adapt CT-CLIP to create a vision-language AI assistant.
We analyze the latent spaces of reports and chest CT volumes generated by CT-CLIP, by computing t-SNE projections of these spaces. Figure 4d shows the distribution according to the number of abnormalities present in a case. Figure 4e depicts the age distribution on the t-SNE projections. The well-clustered t-SNE projections confirm CT-CLIP’s ability to accurately map latent spaces with closely aligned visual and linguistic information. For instance, in both image and text latents, the alignment between age and abnormality distributions indicates that older patients are more prone to abnormalities, which the model effectively identifies. Extended Data Figures 1, 2 provide additional t-SNE projections for each abnormality using volume and report embeddings, respectively.
Surpassing supervised levels with CT-CLIP’s zero-shot multi-abnormality detection
CT-CLIP adopts a zero-shot classification approach for multi-abnormality detection without explicit training on specific abnormality labels. Zero-shot classification refers to the ability of a model to categorize previously unseen classes without being explicitly trained on those specific categories [37]. Instead of relying on predefined labels, CT-CLIP leverages the semantic information embedded in paired radiology reports to infer multiple abnormalities. For instance, a model trained to recognize features like “reticulation,” “consolidation,” or “nodules” in common lung diseases could potentially identify a rare lung condition it was not explicitly trained on by recognizing combinations of these features.
Zero-shot classification with foundation models is typically applied as a multi-class classification task, assigning a single label to each image [37]. However, a medical image may present multiple abnormalities, necessitating a unique approach. Inspired by recent work on 2D chest radiographs [15], we generate positive and negative prompts for each abnormality. The embeddings for the input 3D chest CT volume and the two prompts are projected into a shared latent space, and the contrastive loss between these projections is calculated based on the distance between the image and language spaces (see CT-CLIP for zero-shot multi-abnormality detection section of Methods for more details). This process extracts and identifies the relevant features in both modalities. The model then outputs logits for each prompt, which are interpreted as the likelihood of abnormality presence, as shown in Figure 1e. We conduct a systematic evaluation of CT-CLIP’s zero-shot detection capabilities on our internal validation set from CT-RATE and two external validation sets from Rad-ChestCT and UPMC (Figure 4a).
[…]
To evaluate CT-CLIP’s zero-shot multi-abnormality detection, we benchmark it against CT-Net [20], a state-of-the-art fully supervised model for 3D chest CT volumes. CT-Net requires manual annotation for each predefined abnormality, which demands significant time and expertise [34]. We choose CT-Net as a baseline to demonstrate that CT-CLIP can surpass traditional fully supervised performance without relying on manual annotations or being restricted to predefined classes. Since the pretrained CT-Net model is not available and only a small portion of its dataset is open-source, we train CT-Net on CT-RATE using our abnormality labels for a fair comparison (see benchmarking the CT-CLIP models for multi-abnormality detection section of Methods). Despite using a different dataset for training, the retrained CT-Net’s classification scores are comparable to those reported in the original study [20]. Figure 4a shows the retrained CT-Net’s performance metrics, while Extended Data Figures 3, 4, 5 and Supplementary Figures 2, 6 detail classification scores and confusion matrices.
We compare our zero-shot approach to this supervised baseline and observe significantly higher performance metrics for our approach. On the internal validation set from CT-RATE, the zero-shot CT-CLIP model achieves a mean AUROC that is 0.102 higher and a mean F1 score that is 0.050 higher than the supervised baseline. On the first external validation set, RAD-ChestCT, CT-CLIP outperforms the baseline with a mean AUROC gain of 0.085 and an F1 score improvement of 0.073. Similarly, on the second external validation set from UPMC, CT-CLIP maintains strong performance with a 0.087 increase in mean AUROC and a 0.037 boost in F1 score.Results on the external validation sets demonstrate that the zero-shot CT-CLIP model performs well not only on similar data distributions but also under distribution shifts, indicating its strong generalization to external datasets. A two-sided unpaired permutation test, utilizing 10,000 permutations, highlights a statistically significant improvement $(p<0.05)$ in AUROC as a zero-shot classifier over the supervised (baseline) model, detailed in Supplementary Table 5.
Extended Data Figures 3, 4, 5 provide scores for all abnormalities on the CT-RATE internal, RAD-ChestCT external and UPMC external validation sets, respectively. Supplementary Figures 3, 7 show confusion matrices on the CT-RATE internal and RAD-ChestCT external validation sets, respectively. Notably, on the internal validation set, all abnormalities (18 out of 18) exhibit higher scores in our zero-shot approach compared to the supervised baseline in mean AUROC. In conclusion, CT-CLIP excels by just leveraging existing radiology text reports in electronic health records, thus avoiding the need for expensive manual annotations. Moreover, CT-CLIP’s zero-shot capability offers flexible predictions beyond predefined classes and adapts efficiently to changing needs post-training.
Since the availability of sufficiently large datasets is crucial for machine learning, especially for transformer models, which require more data than convolutional neural networks [40, 41], we explore the impact of dataset size on classification performance. We train CT-CLIP on different portions of the CT-RATE dataset, specifically 9.8% (equivalent to the open-source portion of the RAD-ChestCT dataset [20], currently the largest publicly available chest CT dataset), 20%, 40%, 60%, 80%, and 100%. Figure 4c reports the classification performance across different dataset sizes. Notably, a consistent trend emerges: larger dataset sizes correlate with improved scores. This finding highlights the role of dataset size in enhancing CT-CLIP’s zero-shot classification performance and underscores the value of our extensive CT-RATE dataset for computational research. It also offers a clear avenue for future improvements, suggesting that further expanding the dataset or incorporating more diverse data could lead to even better performance.
Enhancing CT-CLIP’s detection performance with finetuning
To enhance CT-CLIP’s multi-abnormality detection performance, we explore a task-specific finetuning method, ClassFine, inspired by conventional linear probing, by adding a new classification layer to the pretrained vision transformer of CT-CLIP (Figure 3a). Supplementary Figure 1b indicates that training only the final layer may be sufficient for finetuning (see finetuning CT-CLIP for multi-abnormality detection section of Methods). While the ClassFine approach improves detection performance (Figure 4a), it restricts the model to predefined labels as any supervised classification method. Extended Data Figures 3, 4, 5 present abnormality scores, with confusion matrices in Supplementary Figures 5, 9 on the internal and external validation sets, respectively. To address the limitations of ClassFine, we develop the VocabFine approach, a novel open-vocabulary finetuning method inspired by WISE-FT [42] (see finetuning CT-CLIP for multi-abnormality detection section of Methods). VocabFine, like CT-CLIP’s zero-shot inference, is finetuned to improve label prediction accuracy (Figure 3c). Specifically, we use positive and negative prompts for abnormalities, calculate the logits for these volumes with pretrained CT-CLIP, and align these logits with the ground truth labels. While the finetuned model with VocabFine retains the open-vocabulary inference capability of CT-CLIP (Figure 3d), it also consistently outperforms ClassFine (Figure 4a), further emphasizing the value of our open vocabulary finetuning. Abnormality-specific improvements are detailed in Extended Data Figures 3, 4, 5 while Supplementary Figure 1b underscores the importance of training all weights in finetuning. Supplementary Figures 4, 8 provide confusion matrices on the internal and external validation sets, respectively. Notably, both ClassFine and VocabFine are trained solely on CT-RATE; thus, their strong generalization to external sets (RAD-ChestCT and UPMC) demonstrates the robustness of CT-CLIP’s pretrained representations under distribution shifts. This performance advantage across multiple test sets supports the effectiveness of our finetuning strategies. However, despite these promising results, the abnormality detection performance still falls short of clinical deployment thresholds.
Beyond its superior performance, VocabFine retains the open-vocabulary capabilities of the original CT-CLIP model, enabling inference on abnormalities that were not explicitly included during fine-tuning. This makes the model inherently more flexible and clinically useful in real-world scenarios where emerging or rare findings may need to be recognized without retraining. Additionally, unlike ClassFine—which only adds a classification head on top of the vision encoder and is therefore restricted to vision-only downstream tasks—VocabFine preserves and fine-tunes both the vision and text encoders. This dual-modality design allows it to generalize to a wider range of tasks, such as text-to-image retrieval and multimodal question answering, which rely on aligned visual and linguistic representations. In summary, our VocabFine approach not only improves multi-abnormality detection accuracy but also maintains the foundational model’s versatility and adaptability across diverse downstream applications.
Retrieving relevant cases with CT-CLIP
In clinical practice, radiologists frequently refer to past cases with similar findings to guide diagnosis and treatment decisions. This process, known as case retrieval, enhances diagnostic accuracy and builds confidence in decision-making. However, manually searching for comparable cases is time-consuming and inefficient. An automated, robust solution can streamline case retrieval, saving both time and effort.
Beyond zero-shot detection, CT-CLIP serves as a powerful tool for image-to-image retrieval [43], identifying the relevant volumes from a query volume by comparing cosine similarities between the embeddings (Figure 5a). We test on both CT-RATE internal and RAD-ChestCT external validation sets using the vision transformer of CT-CLIP, finetuned models, and the supervised baseline. We utilize the Mean Average Precision at K (MAP@K) metric, where K represents the number of returned images (see benchmarking the CT-CLIP models for CT volume retrieval section of Methods). As K increases, MAP@K scores typically decrease due to less accurate matches. Results, as in Figure 5c, show comparable performance among methods, with all significantly outperforming random retrieval. Although the external validation set from RAD-ChestCT has higher MAP@K scores, the fold increase over random retrieval is lower (three times on the external vs. six times on the internal set), due to differing abnormality distributions. This suggests that examining fold change offers a more accurate comparison. Finetuning does not significantly alter the vision transformer’s weights, as CT-CLIP and the finetuned models perform similarly across MAP@K metrics.
CT-CLIP also functions as a text-to-image retrieval tool [44], evaluating cosine similarities between target volume embeddings and query report embeddings (Figure 5b). Given the limitations of the supervised model and the ClassFine approach in encoding text, we employ CT-CLIP and the VocabFine approach for our experiments. Due to the lack of reports in the RAD-ChestCT dataset, only the internal validation set is used for report-to-volume retrieval. We evaluate performance using Recall@K, where K represents the number of returned images, reflecting the percentage of correctly matched images within the top K returns (see benchmarking the CT-CLIP models for CT volume retrieval section of Methods). While Recall@K scores increase with more returned volumes, considering fold change relative to random performance is more meaningful. Figure 5d shows that CT-CLIP performs significantly higher than the random baseline. However, the Recall@K scores for VocabFine are substantially lower than CT-CLIP, consistent with the expectation that finetuning a foundation model for a specific task might diminish its efficacy for other tasks. The reduced performance with VocabFine in report-to-volume retrieval further indicates significant modifications to the text transformer weights due to finetuning.
CT-CHAT: Developing a multimodal AI assistant with CT-RATE and CT-CLIP
Observing CT-CLIP’s remarkable capabilities for open-vocabulary inference, we build on this foundation to enable an interactive exchange between textual queries and diagnostic image information. To achieve this, we develop CT-CHAT, a multimodal generalist AI assistant specifically designed for 3D chest CT. CT-CHAT integrates both visual and language processing to perform various tasks, making it a versatile tool for clinical use. To train CT-CHAT, we create a visual question-answering (VQA) dataset, comprising over 2.7 million question-answer pairs derived from CT-RATE using Llama 3.1 7B [45] (see creating the dataset for CT-CHAT section of Methods). This dataset consists of several components, including long-answer questions, short-answer questions, multiple-choice questions, and report generation tasks (Extended Data Figure 8), reflecting the diverse types of inquiries encountered in clinical practice.
CT-CHAT’s architecture leverages the strengths of CT-CLIP by using its pretrained vision transformer to encode 3D chest CT volumes. These encodings are then passed to a large language model (LLM) through a multimodal projector, as shown in Figure 6a. We perform an ablation study evaluating CT-CHAT’s performance with 4 different state-of-the-art LLMs, as detailed in Figure 6d. We select 70B Meta Llama 3.1 [45], outperforming others in all key metrics, for the optimal configuration of CT-CHAT.
We create CT-CHAT by finetuning the multimodal projector and Low-Rank Adaptation (LORA) weights [46] for the LLM (see developing the CT-CHAT model section of Methods). Prior to this, we perform a medical concept alignment step, where the LLM is frozen, and the multimodal projector is trained solely to generate radiology reports. This step slightly enhances performance by aligning the multimodal projector with biomedical concepts (Supplementary Table 10b). The ablation study for the volume provision proves CT-CHAT extracts relevant information from the given 3D volumes (Supplementary Table 10a). Although CT-CHAT shows the ability to retrieve information from the given image, it is also expected to hallucinate some pathologies when the input volume is not provided. Thus, we implement a guardrail after the ablation study, with the system prompt (Supplementary Table 9). Supplementary Figure 11 provides an example of the guardrail in case the volume is not provided.
Finally, CT-CHAT generates contextually relevant and clinically accurate responses across a variety of tasks, from simple question-answering to the generation of detailed radiology reports as shown in Extended Data Figure 9. The model is also capable of dynamic task switching, facilitated by special tokens that guide the LLM in generating the appropriate type of response based on the specific task. This flexibility makes CT-CHAT a powerful tool for enhancing radiologists’ workflows, reducing the time and effort required for interpretation, and improving diagnostic accuracy in 3D medical imaging.
Surpassing other multimodal AI assistants with CT-CHAT
Given the absence of generalist models for 3D imaging, we evaluate CT-CHAT against several state-of-the-art open-source vision-language AI models, including LLaVA 1.6 (Mistral 7B and Vicuna 13B) [26], LLaVA-Med [27], and CXR-LLaVA [28]. To ensure a fair comparison, these models, designed for 2D imaging tasks, are tested using Digitally Reconstructed Radiographs (DRRs) of 3D chest CT volumes—a method that ensures accurate representation in a 2D context, as suggested by [47], rather than using random or central 2D slices of a 3D volume (see benchmarking the CT-CHAT model section of Methods). In contrast, CT-CHAT directly processes 3D chest CT, giving it a significant advantage in understanding and interpreting the complex anatomical and pathological information within these scans.
CT-CHAT generates clinically accurate and detailed answers to questions about the given CT image, as shown in Figure 6b. To evaluate the performance, we employ multiple metrics, including BLEU-1, METEOR, ROUGE-L, and CIDEr, which assess the quality of the generated text. Additionally, inspired by previous work [27], we utilize an LLM-based metric, named the Llama score, to evaluate the clinical accuracy of responses, with the models’ outputs compared against ground truths (see benchmarking the CT-CHAT model section of Methods). The results indicate that CT-CHAT, being capable of evaluating the full information of 3D CT volumes, consistently outperforms the other open-source vision-language AI assistants across all evaluated tasks. It shows superior ability in generating long and short answers, handling multiple-choice questions, and producing detailed radiology reports that align closely with clinical expectations, as shown in Figure 6c. Extended Data Figure 6 provides a detailed breakdown, demonstrating that CT-CHAT significantly outperforms others in all tasks.
There are two key reasons why CT-CHAT significantly outperforms 2D-based models: (1) Image: CT-CHAT leverages the 3D spatial information processed by CT-CLIP’s pretrained vision transformer, which captures much richer and more detailed anatomical structures than 2D models—a critical advantage for CT interpretation. (2) Text: CT-CHAT is trained on 3D CT volumes paired with radiology-specific linguistic data from CT-RATE, enabling it to generate highly precise and clinically relevant responses, far surpassing those produced by models trained on general-purpose texts. This powerful combination of CT-CLIP’s capacity to handle complex 3D imaging data and the domain-specific paired image-text data from CT-RATE allows CT-CHAT to push the boundaries of multimodal tasks in 3D medical imaging. In summary, CT-CHAT not only surpasses existing 2D tools but also provides a scalable solution for enhancing chest CT interpretation, underscoring the need for generalist models in 3D medical imaging.
In the absence of existing vision-language baselines for 3D medical imaging, we extend our evaluation of CT-CHAT beyond comparisons with 2D multimodal AI assistants by benchmarking its report generation capabilities against state-of-the-art 3D radiology report generation models—specifically RadFM [29] and CT2Rep [30]. This task is particularly important as it closely reflects the daily workflow of radiologists and provides a more clinically relevant benchmark than VQA. Additionally, the language generation metrics used in our VQA evaluation (such as BLEU and ROUGE) are primarily text-focused and may introduce bias due to differences in model training objectives. They are also limited in assessing the clinical fidelity of responses. Therefore, to robustly evaluate report generation quality, we employ clinical efficacy metrics (precision, recall, and F1 score), which better reflect the diagnostic value of generated reports. We extract these labels using our automated report labeler—the same model used to annotate CT-RATE—and benchmark CT-CHAT using its 8B Meta Llama 3.1 configuration. Notably, even though the RAD-ChestCT dataset lacks text reports, it includes binary abnormality labels, allowing us to evaluate generated reports by applying our text labeler and computing metrics accordingly.
CT-CHAT consistently outperforms both RadFM and CT2Rep on the CT-RATE internal and RAD-ChestCT external validation sets across all three clinical metrics, as shown in Figure 6e. The only exception is precision on the internal set, where RadFM slightly outperforms CT-CHAT. However, this is attributed to RadFM generating fewer positive findings, leading to low false positives—but also to correspondingly low recall and F1 scores—indicating limited clinical utility.
The superior performance of CT-CHAT over CT2Rep is especially noteworthy, given that both models are developed using the CT-RATE dataset. Unlike CT2Rep, which does not leverage pretraining, CT-CHAT benefits from CT-CLIP’s pretrained vision encoder, highlighting the strength of our foundational model. These results reinforce the effectiveness of CT-CLIP in enabling high-quality downstream applications such as clinically meaningful report generation and demonstrate the value of leveraging contrastively learned representations in 3D medical imaging.
Discussion
Radiology reports provide comprehensive descriptions of medical images without adding extra labeling burden, as they are already part of the daily routine for radiologists [48]. Recent studies show that foundational models, trained on paired radiology reports and 2D medical images, perform comparably to fully supervised methods in various tasks [15]. However, these studies have not extended to 3D medical images, primarily because a publicly available 3D medical imaging dataset paired with text reports is absent. Therefore, CT-RATE, which pairs 3D chest CT volumes with radiology reports (Figure 1a), represents a significant advancement as the first of its kind. In this study, we introduce CT-RATE, a novel and comprehensive dataset, which we first utilize to develop CT-CLIP (Figure 1c). To our knowledge, CT-CLIP is the first foundation model in 3D medical imaging utilizing radiology text reports.
CT-CLIP provides a versatile solution for various tasks without requiring task-specific training. Unlike supervised methods, it eliminates the need for explicit labels and enables predictions beyond predefined classes (Figure 1e), making it particularly valuable when learning objectives evolve post-training. CT-CLIP’s zero-shot multi-abnormality detection outperforms the state-of-the-art supervised method on the internal validation set across all metrics (Figure 4a). Importantly, this experiment suggests that explicit labels are not necessary for high performance in 3D medical image interpretation when corresponding radiology reports are available for training. Furthermore, we externally validate CT-CLIP on two independent datasets, Rad-ChestCT and UPMC, from a different country (Figure 3e). The model’s ability to generalize to different data distributions, a primary challenge in deploying medical AI [49], is confirmed as it outperforms the fully supervised method in all metrics in the external validation sets as well (Figure 4a). To establish a fully supervised baseline and fine-tune CT-CLIP, we extract multi-abnormality labels from radiology reports, making them publicly available. Following recent examples, we develop an automatic label extractor by finetuning a language model (Figure 1b) [50]. However, developing such a system requires significant time, expertise, and domain knowledge, as evidenced by our effort in manually annotating a small portion of the reports. CT-CLIP, as a zero-shot classifier, overcomes this bottleneck by detecting abnormalities at a supervised level without the need for an automatic labeler. Finetuning experiments further confirm CT-CLIP’s representational learning capabilities, demonstrating superior performance (Figure 4a). Despite being trained without explicit supervision, CT-CLIP’s vision encoder captures clinically meaningful features that are highly transferable to downstream tasks. These finetuning strategies—one tailored for classification with predefined labels (ClassFine), and the other enabling open-vocabulary inference (VocabFine)—build upon the same pretrained encoder and consistently yield superior results compared to a fully supervised baseline. This indicates that the encoder does not rely on memorizing specific abnormality classes but instead learns a more general and semantically aligned representation of the 3D chest CT volumes. Such flexibility underscores the foundational nature of CT-CLIP and its potential for broad applicability in medical imaging tasks, even beyond those explicitly considered during training. Finally, and indicating a direction for future developments, the effective clustering observed in t-SNE projections (Figure 4d-e) unveils another aspect of CT-CLIP’s clinical potential. The successful clustering of abnormalities (Extended Data Figures 1 and 2) suggests that CT-CLIP could extend its utility beyond multi-abnormality detection to assist in prognosis, disease grading, and the calculation of lung-related parameters in future research.
Analyzing CT-CLIP’s open-vocabulary inference capabilities, we observe significant potential for developing multimodal algorithms that address broader and more diverse clinical applications. To expand these capabilities, we finetune four state-of-the-art LLMs through visual instruction tuning with the CT tokens created by CT-CLIP, resulting in the creation of CT-CHAT. For finetuning, we create and open-source a visual question-answering dataset derived from CT-RATE (Extended Data Figure 8). As a result, CT-CHAT significantly outperforms existing 2D models, underscoring the urgent need for generalist models specialized in 3D medical imaging (Figure 6c). We find the choice of LLM (Supplementary Table 10a) and the use of concept alignment pretraining (Supplementary Table 10b) have minimal impact on CT-CHAT’s accuracy. In contrast, including the volume as input is crucial, particularly in report generation (Supplementary Table 10b). This highlights the value of the CT-RATE dataset and CT-CLIP in developing a generalist AI assistant for 3D medical imaging. Also, we demonstrate that engineering the system prompt of CT-CHAT (Supplementary Table 9) helps mitigate pathology hallucinations when the volume is not provided (Supplementary Figure 11). Furthermore, our model, with new special tokens, is capable of responding in various modes, including long and short answers, multiple-choice, and report generation (Extended Data Figure 9). To further evaluate CT-CHAT’s clinical utility, we benchmark its report generation performance against state-of-the-art 3D medical models RadFM [29] and CT2Rep [30]. Since radiology report generation closely mirrors the daily workflow of radiologists, it serves as a more clinically relevant benchmark than VQA. Instead of relying solely on conventional text similarity metrics—which are limited in assessing the clinical accuracy of generated content—we employ clinical efficacy metrics, using our in-house automated report labeler, which was also used to annotate CT-RATE (see creating the CT-RATE dataset section of Methods for more details). CT-CHAT consistently outperforms both RadFM and CT2Rep on the CT-RATE internal and RAD-ChestCT external validation sets across all clinical metrics (Figure 6e). The only exception is precision on the internal set, where RadFM scores slightly higher due to generating much fewer positive findings. Notably, CT-CHAT’s advantage over CT2Rep—despite both models being trained on CT-RATE—highlights the effectiveness of pretraining with CT-CLIP. These results reinforce the strength of our foundational model, CT-CLIP, in learning clinically meaningful representations for 3D medical imaging and enabling high-quality performance in downstream tasks. That said, despite CT-CHAT setting a new state-of-the-art in report generation for 3D computed tomography, it remains far from being ready for clinical deployment. As with all AI systems, challenges such as robustness, interpretability, regulatory compliance, and prospective validation must be addressed before integration into routine radiology workflows.To help broaden accessibility, we introduce a user-friendly graphical interface (also open-sourced via our GitHub repository) that allows non-experts to interact with CT-CHAT effortlessly. This interface not only makes CT-CHAT a valuable tool for educational and clinical settings but also invites a wider audience to engage with and contribute to the evolution of interactive AI-driven inference in clinical practice.
Despite these advancements, our CT-RATE dataset and the CT-CLIP and CT-CHAT models have certain limitations. First, biases inherent to our single-institution dataset may influence the performance of our self-supervised training. Although we curate a large and diverse set of volume-report pairs, the data originates from a single geographic and clinical setting, which may limit the generalizability of our models to other populations or institutions. We conduct an ablation study to assess the impact of training set size on CT-CLIP’s performance, revealing that performance increases with larger training sets (Figure 4c). This suggests that incorporating data from multiple institutions and more diverse patient populations could further enhance model robustness and generalizability. Second, CT-CLIP depends on the specific terminology used in reports and the choice of prompts during zero-shot detection. Variability in how diseases are described across radiology reports can affect detection accuracy. To assess the importance of prompt engineering, we evaluate zero-shot multi-abnormality detection performance using seven different prompts during inference (Figure 4b). While we select the best-performing prompt from commonly used structures in radiology reports, continued optimization of prompts may further improve performance. Third, the evaluation of CT-CHAT faces limitations due to the lack of pretrained 3D medical multimodal AI assistants for benchmarking. As a result, we compare CT-CHAT against 2D models, which may not fully capture the 3D spatial information inherent in CT volumes. To address this, we utilize digitally reconstructed radiographs (DRRs) generated from our dataset [47]. This approach is suggested by prior work showing that models trained with DRRs can yield comparable results to those trained with 3D chest CT volumes [47]. However, the absence of true 3D benchmarks highlights the need for further development and validation of multimodal AI assistants in 3D medical imaging. Due to the absence of 3D VQA baselines, we also benchmark CT-CHAT’s report generation function against state-of-the-art 3D radiology report generation models. Despite CT-CHAT being trained for VQA, CT-CHAT outperforms these models; however, the quality of its generated reports remains insufficient for clinical use, underscoring the need for further improvements in performance and clinical validation. Finally, our current models are limited to chest CT volumes. To broaden their clinical applicability, future work should focus on extending CT-RATE, CT-CLIP, and CT-CHAT to include other 3D imaging modalities such as MRI and PET scans, as well as other anatomical regions. Expanding to different modalities and body parts would enhance the utility of our models across a wider range of medical imaging tasks and clinical scenarios.
In conclusion, the development of CT-CLIP and CT-CHAT marks a significant advancement in 3D medical imaging, particularly for interpreting chest CT volumes. Enabled by the creation of CT-RATE, the first open-source 3D imaging dataset paired with radiology reports, these models set a benchmark for multimodal AI in 3D medical imaging. The open-source availability of CT-RATE, CT-CLIP, and CT-CHAT is expected to be a valuable asset to the medical AI community, providing a solid foundation for future research. These foundational models also can be deployed in hospitals with minimal finetuning, making them especially useful in emergency settings where radiologists may not always be available, thereby enhancing diagnostic capabilities and knowledge dissemination in 3D medical imaging.
Figure 1: Figure 1: Overview of our dataset and CT-CLIP. a. Data acquisition protocol for the CT-RATE dataset, comprising non-contrast 3D chest CT volumes and corresponding radiology text reports written by expert radiologists. b. Extraction of multi-abnormality labels from radiology reports using a finetuned text encoder. c. Training of our CT-CLIP framework, which pairs chest CT volumes with radiology reports from the CT-RATE dataset through a contrastive learning approach. d. Illustration of the contrastive learning procedure, employed in the correlation matrix during training to ensure precise match recognition and mismatch differentiation. $\text{V}_{\text{i}}$ shows volume embeddings and $\text{R}_{\text{i}}$ shows report embeddings. e. Inference for zero-shot multi-abnormality detection utilizes normalized probabilities derived from softmax-applied contrastive loss for each positive and negative prompt.
Figure 2: Figure 2: Comprehensive analysis of the novel CT-RATE dataset. a. CT-RATE comprises data from 21,304 unique patients, with ages ranging from 18 to 102 years. b. For validation and finetuning purposes, multi-abnormality labels for 18 distinct abnormalities are extracted from the corresponding radiology reports for each CT volume. c. The training process of the CT-CLIP model utilizes the impression and findings sections of radiology reports, which vary in length, along with the chest CT volumes. d. The number of slices in the chest CT volumes within CT-RATE ranges from 100 to 600. e. Spacings in the x-axis, y-axis, and z-axis vary across volumes; all chest CT volumes are preprocessed to ensure consistent spacing before training each model. f. Despite originating from a single institution, CT-RATE includes volumes acquired using scanners from three different manufacturers, ensuring variability.
Figure 3: Figure 3: Finetuning CT-CLIP and the validation strategy. a. Illustration of linear probing finetuning method (ClassFine) for CT-CLIP, where a linear layer is incorporated into the vision encoder. b. ClassFine enables multi-abnormality classification but is limited to the classes predefined during finetuning. c. Illustration of CT-CLIP’s open vocabulary finetuning method (VocabFine) for each abnormality. d. VocabFine allows for open vocabulary abnormality classification even after finetuning, although it is constrained to the prompts provided during finetuning. e. Validation: models are trained on the CT-RATE dataset from Türkiye and then tested on an internal CT-RATE validation set and two external datasets from the U.S.; a detailed analysis of the CT-RATE dataset is presented in Figure 2. f. Comparison: a comprehensive evaluation is performed in multi-abnormality detection across three different cohorts, evaluating CT-CLIP, the two finetuned models, and a fully supervised method.
Figure 4: Figure 4: A detailed evaluation of CT-CLIP a. Multi-abnormality classification performance is evaluated on an internal and two external validation sets using four key metrics. Our CT-CLIP outperforms the fully supervised baseline across all metrics. Performance improves with both finetuning techniques, demonstrating the robustness of the CT-CLIP’s encoder in feature extraction from CT volumes. The scores are color-highlighted to emphasize the performance increase compared to random benchmarks. b.Different prompts are utilized for the zero-shot multi-abnormality detection task with CT-CLIP. The best-performing prompt, highlighting the importance of prompt engineering, is selected for further experiments. c. To explore how the performance scales with dataset size, we train CT-CLIP using different portions of CT-RATE. (9.8% of CT-RATE matches the size of the only currently available open-source chest CT dataset with multi-abnormality labels.) Contour lines denote $\pm$1 standard deviations calculated via bootstrapping. d. t-SNE projections map the distribution of the number of abnormalities present in cases, utilizing the latent values from individual reports and CT volumes in CT-RATE. e. t-SNE projections illustrate the age distribution of patients in the dataset. The clustering shows that CT volumes and reports from older patients tend to correspond with a higher number of reported abnormalities, indicating a correlation between age and disease prevalence.
Figure 5: Figure 5: Utilizing CT-CLIP for retrieval of chest CT volumes a. CT-CLIP can be used for volume-to-volume retrieval utilizing the vision encoder within CT-CLIP. b. CT-CLIP can be used for report-to-volume retrieval utilizing both the vision and text transformers within CT-CLIP. c. Comprehensive evaluation of volume-to-volume retrieval performance on internal and external validation sets, with scores color-highlighted to emphasize improvements over random scores, shows similar retrieval accuracy even for the finetuned models. d. The evaluation of report-to-volume retrieval is limited to the internal validation set due to the absence of paired radiology reports in the external validation set. Notably, finetuning CT-CLIP for detection diminishes its retrieval performance. This reduction in performance is primarily observed in report-to-volume retrieval, indicating that the VocabFine finetuning predominantly impacts on the text transformer component of CT-CLIP.
Figure 6: Figure 6: Overview and detailed evaluation of CT-CHAT a. The architecture of CT-CHAT, a vision-language AI assistant for 3D chest CT volumes, trained by finetuning LoRA weights on a state-of-the-art large language model using CT tokens encoded by the 3D vision encoder of CT-CLIP. b. An illustration showing an example CT-CHAT output, highlighting its superior performance as a vision-language assistant specifically designed for 3D chest CT volumes compared to its 2D counterparts with DRRs provided. c. The question-answering performance of various vision-language assistants is evaluated across five key metrics, with radar charts providing a detailed breakdown of evaluations for each of the four different visual question-answering tasks. d. An ablation study on the accuracy of CT-CHAT using different LLMs, shows that CT-CHAT with the Llama 3.1 model performs slightly better. e. Report generation performance is evaluated using clinical efficacy metrics, calculated by adapting CT-RATE’s automatic report labeler to assess clinical correctness across multiple findings on two test sets.
Online Methods
Ethics statement
Our study is granted with ethical approval from the Clinical Research Ethics Committee at Istanbul Medipol University (E-10840098-772.02-6841, 27/10/2023) for open-sourcing our CT-RATE dataset and associated trained models. To uphold patient privacy, we anonymize all chest CT volumes, associated metadata, and radiology text reports before conducting any computational analysis.
Creating the CT-RATE dataset
We curate a novel dataset, CT-RATE, consisting of chest CT volumes and corresponding radiology text reports from Istanbul Medipol University Mega Hospital. The CT-RATE dataset includes non-contrast chest CT volumes acquired between May 2015 and January 2023, totaling 50,188 reconstructed CT volumes from 25,692 distinct CT experiments conducted on 21,304 unique patients. We divide the cohort into two groups: 20,000 patients are allocated to training and 1,304 for validation. Figure 1a illustrates the dataset accumulation process, and Figure 3e shows the training and validation split.
In clinical practice, CT volumes are reconstructed with various techniques for specific diagnostic requirements. Sharper kernels, for instance, improve resolution for lung abnormalities, whereas smoother kernels are favored for assessing mediastinal issues [32]. For that reason, we incorporate every reconstructed volume from the hospital database into our dataset. This strategy facilitates a thorough evaluation of various abnormalities. The patients range in age from 18 to 102 years, with a mean of 48.8 years and a mode of 40 years. The sex distribution is 41.6% female and 58.4% male. 61.5% of the scans are taken using Philips CT scanners, 30.1% with Siemens scanners, and 8.4% with PNMS (Philips-Neusoft Medical Systems) scanners. The CT volumes are provided in various resolutions: 65.4% in 512 × 512 px, 4.2% in 768 × 768 px, and 30.4% in 1024 × 1024 px. The pixel spacing in the axial (XY) plane ranges from 0.227 to 1.416 mm, with a mean of 0.605 mm, while slice thickness ranges from 0.035 to 6 mm, with a mean of 1.231 mm. The number of slices per volume ranges from 100 to 600, averaging 304.7 with a mode of 255 slices. The radiology reports are structured into four sections: clinical information, technique, findings, and impression. The clinical information section outlines the patient’s symptoms and medical history. The technique section specifies the acquisition protocol, encompassing details such as the field of view. The findings section covers anatomical and pathological observations, whereas the impression offers a diagnosis based on these findings. Figure 2 summarizes the CT-RATE dataset alongside the first external validation set (RAD-ChestCT). Unfortunately, detailed metadata for the second external validation set (UPMC) is not publicly available, as it is a private dataset and cannot be shared; thus, it is not included.
To anonymize the CT-RATE dataset, we first anonymize reports by removing all personal information pertaining to patients and medical professionals. This process involves targeted searches using regular expressions to identify and eliminate date-related information and patient-doctor identifiers. Subsequently, these anonymized reports are translated from Turkish to English using the Google Translate API, and bilingual final-year medical students review and correct all translations. They also thoroughly examine each report to remove any residual personal information and ensure coherence. Only the English versions of these reports are included in our CT-RATE dataset. For the CT volumes, we implement several quality control measures. We exclude volumes that involve contrast material or body parts other than the chest. Only the remaining data is used for developing our models and open-sourced as the CT-RATE dataset. Each chest CT volume’s orientation is standardized by examining its metadata to ensure consistency. Additionally, we remove all identifying information about patients and medical professionals from the metadata. The anonymized metadata is open-sourced as part of the CT-RATE dataset. Extended Data Table 1 provides a detailed explanation of these metadata attributes.
To develop an automated label extractor model, we annotate 1,000 reports by merging findings and impression sections, as some abnormalities are not explicitly mentioned in a single section. This creates a multi-label dataset for each report, identifying 18 abnormalities. Conditions that are not referenced in either section are considered absent. Similar conditions are grouped: left and right mucoid impactions under mucoid impaction, and various lung opacities (lung opacity, density increase, and ground-glass opacities) under lung opacity. Lung and fissural nodules are categorized as lung nodules. The annotated reports are divided into training (80%, 800 reports) and validation (20%, 200 reports) sets. Leveraging this dataset, we finetune the RadBERT-RoBERTa-4m model [35], which was originally trained on over 4 million radiology reports. Then, we apply it to classify the remaining 24,692 reports. These classified reports serve three purposes: training the supervised baseline model, finetuning CT-CLIP, and internal evaluation of the models. Figure 1b illustrates the automated classification process for radiology reports. Supplementary Table 2 shows the precision, recall, and F1 scores of the report classifier.
Our preprocessing approach, inspired by previous work [20], involved several key steps to standardize all chest CT volumes in the dataset. First, each volume is resized to achieve uniform spacing of 0.75 mm in the x-axis and y-axis and 1.5 mm in the z-axis. The CT volumes are then either center-cropped or padded to reach a consistent resolution of $480\times 480\times 240$, ensuring comprehensive inclusion of the patient’s relevant area for the interpretation. Subsequently, we convert each CT volume to Hounsfield Unit (HU) values using intercept and slope values from the metadata. HU values, a measure of radiodensity, are clipped to a range of -1000 to 1000. This range represents the practical lower and upper limits of the HU scale [51]. During training, we normalize these values to a range of -1 to 1. The CT-RATE dataset contains the original raw data and all preprocessing scripts are provided in our GitHub repository.
In addition to radiology reports and metadata, we plan to extend CT-RATE with multi-abnormality segmentation and localization labels for 3D chest CT volumes. To our knowledge, no publicly available dataset currently provides such comprehensive annotations for 3D medical imaging, making this a highly valuable contribution. Given the substantial effort and expert input required, this will be pursued as part of future work. As an initial step, we include organ-level anatomical segmentation labels generated using a state-of-the-art segmentation model [52] (see Supplementary Table 11). These labels are openly available through the CT-RATE repository on Hugging Face and hopefully support tasks such as visual grounding, anatomical localization, and serve as a foundation for future fine-grained abnormality segmentation.
Adding external validation datasets to CT-RATE
For the first external validation set, we utilize the RAD-ChestCT dataset, which includes a cohort from another country [20]. The dataset comprises 3,630 non-contrast chest CT volumes, uniformly reconstructed using a single technique. Half are conducted with Siemens CT, and the other half with GE Medical scanners. The number of slices ranges from 46 to 1277, with an average of 511.6 and a mode of 472. All CT volumes have a resolution of 512 × 512 pixels. The pixel spacing in the axial (XY) plane ranges from 0.189 to 0.977 mm, with an average of 0.692 mm, while the slice thickness varies from 0.125 to 5 mm, with an average of 0.706 mm. To ensure consistent evaluation, we apply identical preprocessing methods to this external dataset as used with our CT-RATE cohort. Since the RAD-ChestCT dataset lacks a Mosaic attenuation label, we exclude it from our external evaluation. Furthermore, while the RAD-ChestCT dataset includes a Calcification label, our dataset distinguishes between Arterial wall calcification and Coronary artery wall calcification. In assessing the model accuracies on the external evaluation set, we use the higher classification probabilities between these two labels for the Calcification prediction.
For the second external validation set, we use the validation split from the MedSyn paper [33], which contains 1,024 chest CT scans and paired radiology reports, each from a different subject. These scans originate from the University of Pittsburgh Medical Center (UPMC) and have been fully de-identified. We employ the same automated report labeler used for CT-RATE to extract 18 abnormality labels from the text reports for evaluation. To ensure consistent and fair comparison, we apply the same preprocessing pipeline to chest CT volumes as used with our CT-RATE cohort.
Developing the CT-CLIP model
CT-CLIP is a novel 3D adaptation of the Contrastive Language-Image Pretraining (CLIP) which is originally designed for natural image-text pairs [37] and has shown excellent performance in 2D medical data [15, 53]. It learns to identify similarities between text and images via contrastive learning (Figure 1d), enabling efficient zero-shot classification. We extend CLIP to work with 3D CT paired with reports. It integrates a 3D encoder, previously used for generating CT [23], as the image encoder and CXR-Bert [50] as the text encoder. Similar to the 2D CLIP, our method extracts latent embeddings from input volumes and radiological reports. We then compute the contrastive loss, which measures the similarity between these embeddings. Training the model to align the 512-dimensional projection layers from text and CT volume embeddings with the identity matrix (as shown in Figure 1c) allows identifying meaningful correlations between textual information and 3D medical images. This process enables zero-shot classification and can be further enhanced through task-specific finetuning for supervised classification.
As a 3D vision transformer, we utilize a component of CT-ViT [23], a network designed to generate low-dimensional tokens from CT volumes and reconstruct CT volumes from these tokens. CT-ViT divides CT volumes into patches measuring $20\times 20\times 10$ voxels in the coronal, sagittal, and axial planes, ensuring uniform spacing. These patches are processed by spatial and causal transformers to produce low-dimensional CT tokens, which are then decoded to reconstruct the CT volumes [23]. For CT-CLIP, we use only the encoder part of this network to generate CT tokens. During the training phase of CT-CLIP, the encoded CT tokens are averaged across the axial plane, then flattened and transformed into a 512-dimensional layer. This creates latent representations for the CT volumes, as shown in Figure 1c. This approach reduces GPU memory requirements significantly through the linear transformation layer for latent representations, while still capturing information from every slice due to the axial attention mechanism in the CT-ViT. The improved memory efficiency enables the use of a larger 3D CT encoder.
For latent representations of radiology reports, we employ the CXR-Bert text encoder [50]. This encoder can process up to 512 tokens, which allows the inclusion of both impression and findings sections from reports. For reports with content of fewer than 512 tokens, we use padding with the $<$PAD$>$ token, where padded tokens are masked during the attention mechanism. Each token is represented in 768 dimensions in the CXR-Bert text encoder. These tokens are then summed across the 512-token span and linearly transformed into a compact 512-dimensional layer, encapsulating the essential textual information from the reports. Figure 1c visually represents this encoding process. Additionally, to assess the impact of including the impression and findings sections, we conduct an ablation study on impression only and findings only training in Supplementary Figure 1a. To highlight the impact of scaling the dataset size on model performance, we conduct an ablation study. This experiment involves training the CT-CLIP model with varying sizes of the CT-RATE dataset: 9.8% (matching the size of the RAD-ChestCT dataset, the only public chest CT dataset with multi-abnormality labels [20]), 20%, 40%, 60%, 80%, and 100%. By systematically altering the sizes of the dataset, our goal is to explore how the accuracy of the model scales with the amount of available training data. Figure 4c illustrates the outcomes of this ablation study.
CT-CLIP for zero-shot multi-abnormality detection
In our CT-CLIP model, we implement a similar approach to CheXzero [15] for zero-shot multi-abnormality classification. This method adapts CLIP’s zero-shot classification technique to handle multiple classes instead of multiple labels for the same class, which is particularly important in medical imaging where images often contain several abnormalities simultaneously. We start by extracting the model’s preliminary output scores, known as logits, using both positive (e.g., “Consolidation is present.”) and negative (e.g., “Consolidation is not present.”) prompts. The prompting method for inference is chosen through prompt engineering experiments. We then apply softmax to these logits to convert them into a likelihood of each abnormality’s presence or absence in a CT volume. Figure 1e illustrates CT-CLIP’s inference strategy for the zero-shot multi-abnormality detection task.
[…]
Finetuning CT-CLIP for multi-abnormality detection
To enhance the performance of the CT-CLIP model in classifying multiple abnormalities, we implement two finetuning approaches. The first approach, called VocabFine, involves finetuning the model while retaining its open vocabulary capabilities. The second approach, adapted from linear probing, is called ClassFine. This method entails training the model with an additional classification layer. We also conduct ablations to investigate the impact of freezing pretrained layers within CT-CLIP during finetuning (Supplementary Figure 1b). Specifically, in addition to training every layer, we finetune only the projection layers of the image and text encoders in the case of VocabFine. For ClassFine, alongside training every layer, we train the model only for the newly added classification layer.
The first approach, VocabFine, is a modification of WISE-FT, a technique for finetuning foundation models [42]. In WISE-FT, the zero-shot model generates logits using classification labels as input text prompts. The model is then finetuned using binary cross-entropy (BCE) [54] to compare these logits with the actual ground truth labels. This process allows the model to learn specific classification labels while maintaining its ability to function with an open vocabulary [42]. Our VocabFine approach adapts this method for the multi-label classification. We generate both positive and negative prompts for each abnormality and then concatenate their respective logits into a unified array of 36 elements (two logits for each of the 18 abnormalities). The ground truth labels are represented as binary arrays (1s and 0s), where $1$ indicates a positive logit if the ground truth is $True$ and vice versa for $False$. For example, if a volume’s label is $True$ for a specific abnormality, it corresponds to $[1,0]$ in the ground truth array; $False$ would be represented as $[0,1]$. The training aligns these ground truth values with the computed logits. To optimize computational efficiency, we divide the 36-element array into smaller segments of 12 elements for each gradient run. Figure 3c depicts the finetuning process in VocabFine. The inference strategies of both CT-CLIP and the finetuned model with Vocabfine are the same (Figure 3d).
Our second finetuning strategy, ClassFine, is achieved by integrating a linear layer into the visual encoder network, a technique inspired by [55]. The architecture of the ClassFine approach is demonstrated in Figure 3a. This modified model undergoes training using ground truth labels. The finetuned model with ClassFine is then evaluated using both internal and external evaluation sets. While this approach offers the potential for improved classification accuracy, it also makes the foundation model more specific for the trained classification task. Since ClassFine is based on fully supervised finetuning, its capabilities are primarily limited to the classes defined during this finetuning phase. Figure 3b visualizes the inference process of ClassFine, highlighting its specific application to the predetermined classes.
Benchmarking the CT-CLIP models for multi-abnormality detection
In assessing our CT-CLIP and finetuned models, we employ the CT-Net model as a fully supervised baseline [20]. Renowned for its performance in a similar CT dataset [20], CT-Net provides a robust comparison point. We train the CT-Net model on our CT-RATE dataset, adapting the chest CT volumes to fit the model’s input requirements by resizing them to $420\times 420\times 402$. Consistent with our models, we focus on the abnormalities present in the CT-RATE dataset (Figure 3d), instead of the 98 and 9 abnormalities explored in the original CT-Net study. For each abnormality, we determine classification thresholds by identifying the left-upper point on the Receiver Operating Characteristic (ROC) curve, using the 10% of each validation set. This analysis is performed across all three evaluation sets—CT-RATE (internal), RAD-ChestCT (external), and UPMC (external)—and for all multi-abnormality detection models: zero-shot, fine-tuned, and fully supervised. Additionally, we compute the Area Under the Receiver Operating Characteristic (AUROC) values for each abnormality.
To provide a comprehensive evaluation, we also calculate accuracy, precision, and weighted F1 scores for each identified pathology. These additional metrics offer more insights into the performance of the models. To further validate the latent spaces generated by CT-CLIP for radiology reports and volumes, we conduct a t-SNE (t-distributed stochastic neighbor embedding) analysis [56]. This analysis enables us to map the positions of abnormalities visually, the distribution of abnormality groups, and other relevant attributes within the t-SNE embeddings (Figures 4d, 4e and Extended Data Figures 1, 2).
For the statistical analysis of our model evaluations, we use the non-parametric bootstrap method to examine the distribution of predictions on internal and external validation sets. This involves 1,000 iterations of random sampling of the volumes with replacement, each using sample sets of size $n$, matching the number of evaluated volumes. We calculate the average metric values for each sample and quantify dispersion by computing the standard deviation of these bootstrap values. For statistical significance, we employ a two-sided paired permutation test with 10,000 permutations to assess differences in performance between the two models across AUROC, weighted F1 score, accuracy, and precision metrics, following previous work [57]. Detailed comparisons are provided in Supplementary Table 5.
Benchmarking the CT-CLIP models for CT volume retrieval
Image retrieval involves identifying the most relevant image from a pool of candidates based on either textual [58] or visual [59] input. With CT-CLIP, trained to associate CT volumes with text reports, volume retrieval becomes possible. This is achieved by computing the cosine similarity among all pairs of volumes and reports, as well as among volume pairs, within a unified embedding space. This embedding space is generated by projecting CT tokens derived from the volumes through the projection layer, Figure 1c. To conduct volume-to-volume retrieval, we calculate the cosine similarity between the latent embeddings of a specific chest CT volume in the validation set and all other abnormal chest CT volumes within the same set. We then rank the volumes based on their cosine similarity scores to identify the most relevant matches for a given volume. This volume-to-volume retrieval process is shown in Figure 5a.
To evaluate this retrieval task’s performance, we adapt the Mean Average Precision at $K$ (MAP@K) metric, typically used in image-to-image retrieval tasks [43]. In this metric, an image is considered relevant if it matches the classes of the target input. Precision is then calculated as the proportion of relevant images within the top $K$ retrieved images. Relevance is determined using the intersection over the union of the abnormality labels marked as $True$ in the ground truth with those in the retrieved chest CT volume. The Average Precision at $K$ (AP@K) is defined as follows:
[…]
Here, $P_{i}$ is the precision at the i-th position, and $R_{i}$ is the relevance at the i-th position. The final class retrieval accuracy is given by:
[…]
Where $n$ is the total number of samples. It is important to note that as $K$ increases, MAP@K may decrease due to the inclusion of less relevant images. We utilize the vision encoders of the supervised baseline, CT-CLIP, and the finetuned models for volume-to-volume retrieval and assess their performance across both internal and external validation sets using MAP@1, MAP@5, MAP@10, and MAP@50 metrics. The Recall@K metric is not applied in image-to-image retrieval due to the inherent limitation that the cosine similarity for identical images would inevitably be the highest, resulting in a consistent Recall@K value of $1$. However, this issue does not arise in text-to-image retrieval since the query is based on text embeddings rather than image embeddings.
For report-to-volume retrieval, we utilize the pretrained text transformer and 3D vision transformer to generate vector embeddings for both radiology reports and CT volumes, which are transformed from the token space to the embedding space with the CT-CLIP’s trained projection layers (Figure 1c). Our goal is to find the CT volume that best matched each query report. To achieve this, we compute the cosine similarity between the latent representations of each report and all the CT volumes. Subsequently, we rank the CT volumes based on their cosine similarity scores to identify and return the most relevant matches. The process of report-to-volume retrieval is depicted in Figure 5b.
For the evaluation of this task, we employ the Recall@K metric, where $K$ represents the number of returned images. This method serves as a standard evaluation tool in text-to-image retrieval tasks [44]. We assess four Recall@K metrics: Recall@5, Recall@10, Recall@50, and Recall@100. These metrics measure the frequency of the target image appearing among the top 5, 10, 50, and 100 retrieved images, respectively. It is worth noting that the external validation set lacks text reports. Therefore, we restrict the report-to-volume retrieval evaluation to the internal validation set. We assess both CT-CLIP and the finetuned model with VocabFine for this task. However, since the supervised model and ClassFine do not generate latent embeddings for reports, they are not utilized in the report-to-volume retrieval.
Creating the dataset for CT-CHAT
To train CT-CHAT, we curate a comprehensive visual question-answering (VQA) dataset derived from free-form reports and chest CT volumes, using the same train-validation split as CT-RATE. The dataset is composed of four main components, with the following details for the training set.
The first component is the long-answer questions, which are further divided into four subcategories: conversation-based questions, description questions, free-response questions, and text-only questions. The conversation, description, and free-response questions are generated from radiology reports, while the text-only questions are created using 48,343 articles related to chest CT, retrieved from the PubMed Central-OA (www.ncbi.nlm.nih.gov/pmc/tools/openftlist). In total, the long-answer component of the training set includes 281,689 conversations, with a total of 518,053 conversation questions, 138,104 description questions, 138,061 free-response questions, and 250,760 text-only questions. All questions are generated using the Llama 3.1 8B large language model [45], following the methodology described in [14]. The scripts for generating the question-answer pairs are available in the GitHub repository, and the prompts are provided in Supplementary Table 7. The second component consists of multiple-choice questions, also generated by Llama 3.1 8B [45], following the same approach described in [14]. This component of the training set includes a total of 46,911 multiple-choice conversations, comprising 138,608 questions in total. The third component includes short-answer questions derived from radiology reports in the CT-RATE dataset using a rule-based approach, as described in [60]. These questions cover categories such as location, presence, abnormality, and disorder. A total of 1,417,828 short-answer questions are sampled to create 318,710 short-answer conversations in the training set. The fourth component involves radiology report generation questions, totaling 47,149 conversations in the training set, where each conversation contains a single report generation question along with the corresponding generated report. The report generation questions are randomly selected from a predefined set of questions (Supplementary Table 6).
The validation set is curated similarly to the training set, using the free-form reports from the CT-RATE validation set. The validation dataset includes 33,326 conversation questions, 8,886 description questions, 8,888 free-response questions, 12,013 short-answer questions, 8,915 multiple-choice questions, and 3,039 report-generation questions. Extended Data Figure 7 shows example entries from the described dataset. Since the VQA dataset is generated by a large language model, we carefully select only accurately generated and properly formatted data—that is, data that can be successfully parsed as JSON—for training our models. Additionally, we limit the initial number of generated questions to maintain a reasonable training duration. Our VQA dataset is made publicly available alongside CT-RATE.
Developing the CT-CHAT model
Adapted from the LLaVa framework [61], CT-CHAT integrates the frozen CT-CLIP vision encoder, a multimodal projector module, and a large language model (LLM) to enhance multimodal processing capabilities. The visual encoder compresses 3D chest CT volumes into a low-dimensional feature space. The multimodal projector module then transforms these visual features into the LLM’s embedding space, aligning it with text tokens. The LLM processes these visual-language tokens as input to generate contextual responses. Figure 6a summarizes the architecture of the CT-CHAT model.
In our implementation, we evaluate and compare four different LLMs. For the primary configuration of CT-CHAT, we use the 70B Meta Llama 3.1 [45], one of the state-of-the-art open-source LLMs. This model consists of 80 transformer layers, 64 attention heads, an embedding dimension of 8,192, an FFN dimension of 28,672, and supports a context size of up to 128k tokens. For a smaller model, we utilize the 8B Meta Llama 3.1 [45], which includes 32 transformer layers, 32 attention heads, an embedding dimension of 4,096, an FFN dimension of 14,336, and a context size of up to 128k tokens. Additionally, for comparison, we implement two LLMs used in the official LLaVA 1.6 [26]: Vicuna 13B and Mistral 7B. Figure 6d shows the scores for each of these LLMs. To dynamically adjust the model’s output mode based on the specific task, we add four special tokens to the tokenizers of each model: <long_answer>, <short_answer>, <report_generation>, and <multiple_choice>. These tokens are selected for different tasks during training and inference. The multimodal projector, based on the approach described in [14], consists of an attention-pooling layer followed by a 2-layer multi-layer perceptron (MLP). The attention-pooling mechanism utilizes 256 learned latent queries in conjunction with multi-headed cross-attention to compress the final layer’s features from the encoder backbone into a fixed-length sequence of image tokens. This approach optimizes both training and inference, ensuring that the sequence length remains within the LLM’s context window, and making the model focus more on the important features with the attention mechanism. The MLP that follows is adapted from LLava 1.6 [26] and includes a single hidden layer activated by GeLU, transforming the image tokens to match the LLM’s embedding dimension.
Training CT-CHAT involves two steps. First, in the medical concept alignment step, the LLM is frozen, and the multimodal projector is trained with the input “Please provide the radiology report for the following chest CT volume <image>” and the output report. This step improves subsequent finetuning performance and helps the multimodal projector to be aligned in biomedical concepts [27]. We conduct an ablation study of this step to show its effect in Supplementary Table 10b. In the second step, medical instruction tuning, the model is trained with the CT-VQA dataset that includes long answer, short answer, multiple choice, and report generation question-answer pairs. During this step, the multimodal projector, along with Low-Rank Adaptation (LORA) weights [46] for the LLM with rank 128 and alpha 256, is finetuned. <image> tokens are added to indicate the embedded image input to the model, except for text-only questions and answers. The four special tokens are added to the respective question-answer pairs in the dataset, enabling the model to learn different output modes. These tokens are later used during inference to select different modes of the model. The hyperparameters for each model for pretraining and finetuning are provided in Supplementary Table 4a, 4b.
Benchmarking the CT-CHAT model
Given the lack of pretrained VQA models for 3D imaging, we conduct a comparative analysis of CT-CHAT against four 2D state-of-the-art open-source VQA model: LLaVA 1.6 (Mistral 7B) [26], LLaVA 1.6 (Vicuna 13B) [26], LLaVA-Med [27], and CXR-LLaVA [28], using benchmarking framework established in [14]. To address the challenge of representing 3D information using their 2D encoders and to ensure a fair comparison, we utilize Digitally Reconstructed Radiographs (DRRs) derived from CT-RATE, as suggested by [47]. This approach is chosen over the selection of random or central CT slices, which can not adequately capture all relevant pathologies. The use of DRRs is justified, as [47] demonstrates that models trained with DRRs show comparable results to those trained with 3D chest CTs.
To evaluate the baseline models and CT-CHAT models with different LLMs, we employ four NLP metrics: BLEU-1 [62], METEOR [63], ROUGE-L [64], and CIDEr [65], along with the Llama score, which measures the clinical accuracy of the generated responses compared to the ground truth, using the Llama 3.1 70B model [45]. A similar LLM-based clinical accuracy measurement approach is used in Llava-Med with the GPT-4 API [27]. Supplementary Table 8 presents the system prompt for the Llama 3.1 model, which is used to score the generated answers. We calculate scores task specifically: long answer, short answer, multiple choice, and report generation. Figure 6c illustrates mean and task-specific scores for each model. To visualize the clinical accuracy of the models, we categorize the Llama scores for every question in each task into three ranges: 0-3 for low clinical accuracy, 4-6 for average clinical accuracy, and 7-10 for high clinical accuracy, and create Sankey plots within these ranges in Supplementary Figure 10.
In addition to benchmarking CT-CHAT on VQA tasks, we further evaluate its performance in radiology report generation—a task more reflective of routine radiological workflows—by comparing it with two state-of-the-art 3D report generation models: RadFM [29] and CT2Rep [30]. While VQA metrics such as BLEU, ROUGE, and the Llama score help assess linguistic and clinical quality, they may not fully capture diagnostic utility. Therefore, for this comparison, we employ clinical efficacy metrics—precision, recall, and F1 score—using our automated report labeler, which extracts structured abnormality labels from generated reports. Evaluation is conducted on both the internal CT-RATE validation set and the external RAD-ChestCT dataset. For RAD-ChestCT, which lacks text reports, we leverage its binary abnormality annotations to enable indirect evaluation through our labeler. Importantly, we use the original pretrained weights for both RadFM and CT2Rep without retraining.
CT2Rep is trained on the same CT-RATE dataset as CT-CHAT and uses the same architecture for its vision encoder. However, unlike CT-CHAT, which benefits from a contrastively pretrained CT-CLIP encoder, CT2Rep is trained from scratch without any pretraining. Additionally, CT2Rep is trained specifically for report generation, while CT-CHAT is trained more broadly across multiple VQA tasks—including classification, question-answering, and report generation. Despite this, CT-CHAT consistently outperforms CT2Rep across all clinical efficacy metrics. This performance gap highlights the substantial value of contrastive pretraining with CT-CLIP in capturing clinically meaningful representations of 3D CT volumes, which can then be effectively transferred to downstream tasks such as report generation—even when the model is not trained exclusively for that task.
To evaluate CT-CHAT’s ability to retrieve information from chest CT, we conduct an ablation study for the image provision, as shown in Supplementary Table 10a. Since the model is expected to hallucinate pathologies when the input volume is not provided, we implement a guardrail for prevention with the system prompt in Supplementary Table 9. Supplementary Figure 11 provides an example of the guardrail.
Computational hardware and software
Experiments for CT-CLIP are conducted using Python 3.11.5 with PyTorch 2.0.1, CUDA 11.7, and other libraries like SciPy 1.11.1, Torchvision 0.15.2, Scikit-learn 1.3.0, Pandas 2.0.3, NumPy 1.24.4, Transformers 4.30.1, and Accelerate 0.21.1. For CT-CHAT, Python 3.12.4 is used with PyTorch v2.4.0, CUDA v12.4, SciPy v1.14.0, Torchvision v0.19.0, Scikit-learn v1.2.2, Pandas v2.2.2, and NumPy v1.26.4. We use four 80GB A100 GPUs, 500GB RAM, and 24 CPUs, mainly for CT-CLIP training with Distributed Data-Parallel and Fully Sharded Data Parallelism. Finetuning of CT-CLIP and other experiments is performed on a single A100 GPU. CT-CHAT with the Llama 3.1 70B model is trained on four A100 GPUs using DeepSpeed ZeRO-3, while other CT-CHAT models use two GPUs. Hyperparameters are in Supplementary Tables 3, 4. All software details are available on our respective GitHub repositories.
Data availability
Our CT-RATE dataset, comprising paired chest CT volumes and radiology text reports, is publicly accessible via https://huggingface.co/datasets/ibrahimhamamci/CT-RATE. The external validation set, RAD-ChestCT, is also openly accessible through https://zenodo.org/records/6406114.
Code availability
Our trained models and codebase are publicly available at https://github.com/ibrahimethemhamamci/CT-CLIP and https://github.com/ibrahimethemhamamci/CT-CHAT for further research.
Author contributions
I.E.H., S.E., and B.M. designed the study. I.E.H. and S.E. were responsible for data collection, data analysis, model construction, model validation, figure preparation, and manuscript writing. F.A., A.G.S., S.N.E., and I.D. managed data anonymization and annotation. M.F.D. contributed to model construction. C.W., W.D., and K.B. contributed to the external evaluation. O.F.D. contributed to the graphical user interface and manuscript writing. M.X. and T.A. contributed to the creation of anatomical segmentation labels. B.W., E.S., M.S., E.B.E., A.A., A.S., B.L., C.B., and M.K.O. provided knowledge support. B.M. supervised the all study. All authors reviewed the manuscript and approved the final version.
Acknowledgments
We extend our gratitude to the Helmut Horten Foundation for their invaluable support, which made this work possible. Additionally, we would like to express our sincere appreciation to Istanbul Medipol University Mega Hospital for their support and provision of data. C.B. additionally receives research support from the Promedica Foundation, Chur, Switzerland. C.W., W.D., and K.B. receive funding from R01HL141813 (An Integrative Radiogenomic Approach to Design Genetically-Informed Image Biomarkers for Characterizing COPD), the Hariri Institute for Computing’s Junior Faculty Fellows Program, and the Hariri Institute Focused Research Program. Figures 1, 3, 5, and 6 are created with BioRender.com.